OpenStudy (anonymous):

Find the distance between the points A(10, –9) and B(8, 7). Round your answer to the nearest hundredth.

OpenStudy (texaschic101):

there is a formula for this... d = sq rt (x2 - x1)^2 + (y2 - y1)^2 (10,-9) x1 = 10 and y1 = -9 (8,7) x2 = 8 and y2 = 7 now we sub d = sq rt (8 - 10)^2 + (7 - (-9)^2 d = sq rt (-2^2) + (16^2) d = sq rt (4 + 256) d = sq rt 260 d = 16.1245 rounded = 16.12 read this... http://www.mathwarehouse.com/algebra/distance_formula/index.php